OpenStudy (anonymous):

Will someone please explain how this works to me? I'm going to post the actual question as a comment, need to enter it in equation form. Deals with fractional exponents and the like.

OpenStudy (anonymous):

\[256^{3/4}\] becomes \[\sqrt[4]{16,777,216}\] I need to know how to find the answer. My calculator has a button to figure it out, but that feels like cheating. =)

OpenStudy (mertsj):

Find the 4th root of 256 first and then do the power. Keeps the numbers smaller.

OpenStudy (mertsj):

In other words think of it like this: \[256^{\frac{3}{4}}=(256^{\frac{1}{4}})^3\]

OpenStudy (anonymous):

Okay, but how do I do it? My book isn't very good at explaining things...it just kinda gives an equation with letters and then replaces it with familiar numbers, and expects ya to know how to manage the big ones. Most annoying!

OpenStudy (mertsj):

Find a number, by trial and error, that if you raise it to the 4th power you get 256.

OpenStudy (mertsj):

It can't be a very big number because 10^4 is 10000 and that is way too big.

OpenStudy (anonymous):

It's four?

OpenStudy (mertsj):

yes. Now you know the 4th root of 256 is 4 so you can finish the problem which says to raise that to the third power.

OpenStudy (anonymous):

Okay! That makes much more sense! Thanks!

OpenStudy (mertsj):

yw